OpenStudy (kainui):

gr8 equation proove it

OpenStudy (kainui):

\[\huge \prod_{n=1}^\infty e^n = \frac{1}{e^{1/12}}\]

OpenStudy (unklerhaukus):

\[\prod_{n=1}^\infty e^n\\ =e\cdot \prod_{n=2}^\infty e^n\\ =e\cdot e^2\cdot \prod_{n=3}^\infty e^n\\ =e^{1+2}\prod_{n=2}^\infty e^n\\ =e^{1+2+\dots}\\ = e^{\displaystyle\sum_{n=1}^\infty n}\\ =e^{-1/12}\]

OpenStudy (unklerhaukus):

zat it?

OpenStudy (kainui):

beutufel. medal+fan

rebeccaxhawaii (rebeccaxhawaii):

*admires brain* @UnkleRhaukus

OpenStudy (unklerhaukus):

(a little bit neater) \[\prod_{n=1}^\infty e^n\\ =e\cdot \prod_{n=2}^\infty e^n\\ =e\cdot e^2\cdot \prod_{n=3}^\infty e^n\\ =e^{1+2+3}\cdot\prod_{n=4}^\infty e^n\\ =e^{1+2+3+4+\cdots}\\ = e^{\sum_{n=1}^\infty n}\\ =e^{-1/12}\]

OpenStudy (astrophysics):

Rofl, that seems so simple in hindsight

OpenStudy (unklerhaukus):

the -1/12 gave it away

OpenStudy (astrophysics):

Bravo

OpenStudy (unklerhaukus):

We have shown that 1 > (2.718...) x (2.718... x 2.718...) x (2.718... x 2.718... x 2.718...) x ...

OpenStudy (kainui):

i see no prblems here B)

OpenStudy (kainui):

while we're at it: \[i=e^{(\pi/2)*i}\] therefore you can get this complex number by only raising a real number to an exponent infinitely...! :D \[i =e^{(\pi/2)e^{(\pi/2)e^{(\pi/2)e^{\cdots}}}}\]